### v1

# KVmix: Gradient-Based Layer Importance-Aware Mixed-Precision Quantization for KV Cache

Abstract
The high memory demands of the Key-Value (KV) Cache during the inference of Large Language Models (LLMs) severely restrict their deployment in resource-constrained platforms. Quantization can effectively alleviate the memory pressure caused by KV Cache. However, existing methods either rely on static one-size-fits-all precision allocation or fail to dynamically prioritize critical KV in long-context tasks, forcing memory-accuracy-throughput tradeoffs. In this work, we propose a novel mixed-precision quantization method for KV Cache named KVmix. KVmix leverages gradient-based importance analysis to evaluate how individual Key and Value projection matrices affect the model loss, enabling layer-specific bit-width allocation for mix-precision quantization. It dynamically prioritizes higher precision for important layers while aggressively quantizing less influential ones, achieving a tunable balance between accuracy and efficiency. KVmix also introduces a dynamic long-context optimization strategy that adaptively keeps full-precision KV pairs for recent pivotal tokens and compresses older ones, achieving high-quality sequence generation with low memory usage. Additionally, KVmix provides efficient low-bit quantization and CUDA kernels to optimize computational overhead. On LLMs such as Llama and Mistral, KVmix achieves near-lossless inference performance with extremely low quantization configuration (Key 2.19bit Value 2.38bit), while delivering a remarkable 4.9× memory compression and a 5.3× speedup in inference throughput.
1 Introduction
Large Language Models (LLMs) [1], such as GPT [2], Llama [3], and their derivatives, have significantly advanced the field of Natural Language Processing (NLP). These models exhibit outstanding performance [4, 5] across a diverse array of tasks, including text generation, question answering, and machine translation. The Key-Value (KV) Cache plays an essential role in the autoregressive decoding process of LLMs. The KV Cache substantially reduces redundant computations in the attention mechanism by storing Key and Value states from preceding time steps for subsequent token generation, thereby significantly accelerating inference efficiency [6]. Nevertheless, as sequence lengths grow, the memory footprint of the KV Cache increases linearly, presenting a formidable challenge to hardware resources. For example, a model with 70 billion parameters may demand over 50GB of memory to maintain the KV Cache when generating a sequence of 20k tokens—a requirement that far exceeds the memory capacity of typical GPUs. In scenarios involving multiple concurrent requests, the KV Cache for each request cannot be shared due to its dependence on unique preceding prompts. Although model parameters can be reused, memory quickly becomes saturated due to the extensive KV Cache demands. Once memory is depleted, data is offloaded to system memory or even disks, resulting in frequent High Bandwidth Memory (HBM) exchanges with system memory. This process causes latency to surge exponentially, leading to catastrophic performance degradation.
The characteristics of the KV Cache outlined above severely restrict the practical deployment and inference efficiency of LLMs in resource-constrained environments, highlighting the pressing need for efficient memory optimization strategies [7, 8]. In recent years, research tackling this issue has predominantly focused on reducing the memory overhead of the KV Cache through quantization and sparsification techniques [9, 10, 11]. Quantization methods, in particular, have gained widespread adoption in industry, significantly contributing to the scalability and accessibility of large-scale models [12]. By quantizing the KV Cache, memory usage can be markedly reduced—for instance, transitioning from FP16 to INT8 halves memory requirements while accelerating memory access and improving throughput. Existing quantization methods have demonstrated impressive model performance even at very low bit-widths. However, these methods either rely on static one-size-fits-all precision allocation schemes [13, 14], lacking flexibility and performance-aware adaptation capabilities, or suffer from the high computational overhead of dynamic quantization while failing to adaptively prioritize critical KV pairs in long-context tasks [15, 16]. Therefore, they are forced to make suboptimal trade-offs among memory usage, model accuracy, and computational throughput.
To address these problems, this paper proposes KVmix, a novel mixed-precision quantization method for KV Cache. KVmix evaluates KV importance using gradient norms of Key and Value projection weights relative to the model’s loss function, enabling layer-specific mixed-precision quantization with flexible bit-width allocation. It proposes a dynamic long-context optimization strategy, keeping full-precision KV pairs for recent pivotal tokens while compressing older pairs. Additionally, KVmix offers low-bit compression strategies and efficient CUDA implementations. KVmix maximizes KV Cache compression and model throughput while maintaining controlled precision. Our specific contributions are as follows:
  • We propose a novel layer importance-aware mixed-precision quantization method. This method assesses the importance of KVs at each layer by computing the $L2$ gradient norms of Key and Value projection weights with respect to the model’s loss function. Based thereon, it independently applies mixed-precision quantization to different layers, allocating higher bit-widths to critical layers and lower to less influential ones. Therefore, it provides the flexibility to balance between accuracy and resource efficiency across diverse inference scenarios.
  • We propose a dynamic pivotal context selection strategy to optimize long-context tasks. According to the KV importance analysis, it adaptively keeps full-precision KV pairs for recent pivotal tokens while aggressively compressing older pairs. This strategy ensures high-quality sequence generation in long-context inference scenarios while dynamically reducing the number of full-precision KV pairs for better memory optimization.
  • We design efficient CUDA implementations and a high-compression 3-bit quantization method for KVmix. Extensive experimental results show that KVmix achieves nearly lossless model accuracy across multiple LLMs and datasets, as well as a 4.9$\times$ reduction in runtime memory usage and a 5.3$\times$ speedup in inference efficiency, outperforming prior state-of-the-art (SOTA) quantization methods for KV Cache.
2 Related Work and Motivation
2.1 Related Work
To mitigate the memory challenges associated with KV Cache, researchers have developed a variety of optimization approaches, primarily centered on compression techniques and dynamic memory management [17, 18]. We mainly discuss the KV Cache compression techniques related to this work. Existing compression approaches encompass quantization, sparsification, and KV Cache sharing [19, 20, 21].
Extensive research has focused on reducing the memory overhead of KV Cache through quantization. For example, KIVI [13] introduced a 2-bit asymmetric quantization technique, employing per-channel quantization for Keys and per-token quantization for Values. This method substantially improves model throughput while preserving accuracy at low bit-widths. KVQuant [22] proposed a non-uniform quantization strategy, integrating pre-RoPE per-channel Key quantization with per-token Value quantization. It employs offline calibration to manage outliers, achieving robust performance in long-context inference scenarios. QAQ [15] developed a dynamic mixed-precision quantization method that calculates quantization bits for Keys and Values online and optimizes the trade-off between accuracy and compression ratio by predicting attention scores. Atom [23] investigated a mixed-precision scheme involving 4-bit and 8-bit activations, dynamically quantizing activations to adapt to input distributions. SKVQ [16] implemented a sliding window quantization method, maintaining high precision for recent KV pairs while reducing quantization errors through channel reordering. IntactKV [14] simplified the quantization process by preserving critical tokens (e.g., initial tokens) with full precision, thereby sustaining generation quality. WKVQuant [24] further compressed the KV Cache by incorporating weight quantization. Numerous other studies have also made significant contributions to KV Cache quantization [25, 26, 27]. These efforts generally aim to quantize as many KVs as possible to the lowest feasible bit-width. While methods such as KVQuant, QAQ, and Atom have introduced mixed-precision quantization for KV, their primary focus is on addressing outliers that are sensitive to precision changes. In contrast, this work proposes a more flexible quantization framework, enabling users to balance average quantization bit-width and model accuracy during deployment based on their specific needs.
In addition to quantization, sparsification techniques have been widely used to optimize KV Cache. These approaches typically leverage importance assessments derived from the attention weight to dynamically adjust KV storage requirements. H2O [28] introduced a resampling-based strategy that retains KV entries with high attention scores for efficient memory compression. Scissorhands [29] capitalized on the "importance persistence" hypothesis to dynamically evict low-importance KVs, which further improves sparsification efficiency. SnapKV [30] employed predictive modeling of future attention patterns to preemptively discard irrelevant KV entries, while PyramidInfer [31] extended H2O by recognizing that the number of critical KV entries diminishes with increasing model depth, progressively increasing sparsification across layers. We observe that evicting KV from shallow layers propagates the effect to deeper layers through hidden states. Unlike PyramidInfer, we use the contribution of KV to the model output to evaluate the importance of different layers. Our method is orthogonal to existing sparsification methods and can also be used as a guideline for KV sparsification to achieve more accurate KV eviction.
Figure 1: Projection matrix weights of K and V across different layers for the Llama 2-7B model. The heatmap visualizes the projection weights by sampling every sixth element. Additionally, the weight values along the axes have been normalized. "Norm" represents the L2 norm of the weight matrix for each layer, while "Range" indicates the range of values within each layer’s weight matrix.
2.2 Motivation
Current KV Cache quantization methods depend on fixed quantization strategies, which uniformly apply the same bit-width to the Keys or Values across all layers of a model. These methods fail to account for the differing contributions of Keys and Values from various layers to the final output. For each layer of the KV Cache, the computation process is outlined as follows: at time step $t$ and for the $i$-th layer, the input comprises hidden states (denoted as $H_{i-1,t}$), which are supplied by the output of the preceding layer. These hidden states are then utilized to compute the K and V for the current token: $K_{i,t}=W_{k_{i}}\cdot H_{i-1,t}$ and $V_{i,t}=W_{v_{i}}\cdot H_{i-1,t}$, where $W_{k_{i}}$ and $W_{v_{i}}$ represent the projection weights for the K and V, respectively, at the $i$-th layer. Subsequently, the computed $K_{i,t}$ and $V_{i,t}$ are concatenated with the previously stored KV Cache, yielding the complete K and V sequences up to time step $t$: $K_{i,1:t}=[K_{i,1},K_{i,2},\dots,K_{i,t}]$ and $V_{i,1:t}=[V_{i,1},V_{i,2},\dots,V_{i,t}]$. This computation process reveals that the projection weights $W_{k_{i}}$ and $W_{v_{i}}$ determine how the model extracts K and V information from the hidden states. These weights directly influence the quality of the KV generated within the attention mechanism, thereby forming the foundation of the layer’s computation. As a result, the weights of the KV affect the layer’s contribution to the model’s overall output.
To substantiate this analysis, Fig. 1 provides heatmaps of the projection weights $W_{k}$ and $W_{v}$ for the Llama 2-7B model [32], with additional heatmaps for other models included in Appendix A. Examination of these heatmaps yields two critical insights: ① There is considerable variation in the weight values of KV across different layers of the model.Within the same layer, the Key and Value weights display distinct distributional patterns. These observations highlight the fact that the influence of Keys and Values from different layers on the model’s final output is not uniform.
3 Methodology
3.1 KV Importance Analysis
In the self-attention mechanism, the Key and Value are computed by applying linear transformations to the hidden states of the previous layer using KV projection weight matrices and are subsequently combined with the Query to compute the attention output. Consequently, the magnitude of KV projection weights alone is insufficient to measure the importance of the Key and Value of each layer, and a more precise evaluation metric is needed. To address this problem, we propose the KVmix profiler, a gradient-based method that quantifies the contribution of each layer’s Key and Value to the model’s output, enabling a judicious mixed-precision quantization strategy. Specifically, we compute the $L2$ norm of the gradients of the model’s loss function with respect to the Key and Value projection weight matrices for each model layer, and evaluate the importance of the Key and Value components based on these $L2$ norm values. KVmix profiler captures the dynamic sensitivity of these parameters during the model inference process, and provides a layer-specific importance metric to support efficient mixed-precision quantization in subsequent inference stages.
The implementation of KVmix profiler consists of the following three key steps: ① Data preparation and forward propagation. A full-precision model is loaded, and multiple prompts are randomly sampled from a target dataset to serve as input data. These prompts are tokenized into input tensors using the tokenizer. Leveraging the autoregressive property of LLMs, each input tensor is shifted left by one position to be used as the corresponding label tensor for computing the model’s loss function. Subsequently, the loss value for each input is determined through forward propagation. ② Gradient calculation and importance assessment. For the $i$-layer of the model, the gradients of the losses with respect to the projection weights of the Key ($W_{k_{i}}$) and Value ($W_{v_{i}}$) are computed independently. This process begins with backpropagation to calculate the gradients, i.e., $\nabla_{W_{k_{i}}}\mathcal{L}$ and $\nabla_{W_{v_{i}}}\mathcal{L}$, where $\mathcal{L}$ represents the loss function. The magnitude of these gradients is then evaluated using the $L2$ norm, i.e., $\|\nabla_{W_{k_{i}}}\mathcal{L}\|_{2}$ and $\|\nabla_{W_{v_{i}}}\mathcal{L}\|_{2}$. The importance scores of the Keys and Values for each layer can be expressed as:
$s_{k_{i}}=\|\nabla_{W_{k_{i}}}\mathcal{L}\|_{2},s_{v_{i}}=\|\nabla_{W_{v_{i}}}\mathcal{L}\|_{2}$ (1)
A larger $s_{k_{i}}$ or $s_{v_{i}}$ signifies a greater impact of that $i$-th layer’s Key or Value on model’s output. To enhance assessment reliability, the gradient norms can be averaged across multiple prompts ($p$), yielding an average importance score for each layer’s Key and Value:
$\bar{s}_{k_{i}}=\frac{1}{P}\sum_{p=1}^{P}s_{k_{i}}^{(p)},\bar{s}_{v_{i}}=\frac{1}{P}\sum_{p=1}^{P}s_{v_{i}}^{(p)}$ (2)
Using the importance scores, we classify the importance of the Key and Value components across all model layers. The top 20% of the layers of $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ are quantized with high-bit representations (e.g., 3-bit or 4-bit), while the remaining 80% of layers adopt more aggressive low-bit quantization (e.g., 2-bit). This 20%-80% split is not fixed and can be dynamically adjusted according to the requirements to balance the trade-off between model accuracy and memory usage. Increasing the proportion of low-bit quantization layers can further reduce the memory footprint of the KV Cache, but may sacrifice some accuracy. ③Model Configuration and Inference. The KV quantization configuration results derived from the above steps are incorporated into the model configuration, enabling the quantized model to be used for inference. The detailed workflow is depicted in Fig. 2, with the algorithmic procedure outlined in Algorithm B of Appendix B. The profiling is performed offline and therefore does not affect the inference efficiency. In addition, it is once-and-for-all for the inference of the model.
Figure 2: The overview of KVmix profiler.
3.2 Asymmetric Low-Bit Quantization
Asymmetric Quantization Strategy. We use per-channel and per-token grouping quantization methods for Key and Value, respectively. The KV Cache has the shape $[B,nh,T,D]$, where $B$ is the batch size, $nh$ is the number of attention heads, $T$ is the token sequence length, and $D$ is the head dimension. When the Key is quantized per channel ($D$), the tensor is reshaped to $[B\times nh\times D,T]$, with each group comprising all tokens of a single channel. This approach is inspired by the distributional properties of the Key Cache that exhibit significant outliers in the channel dimension, i.e., certain channels exhibit significantly large magnitude values. Per-channel quantization isolates errors within each channel and prevents outliers from affecting other channels. When Value is quantized per token, the shape of the tensor is preserved, and each group contains all channels of a single token. Unlike the Key Cache, the Value Cache has no pronounced outliers, but plays a critical role in computing the attention output. Per-token quantization confines errors to individual tokens, preserving the integrity of other important tokens. This asymmetric quantization strategy effectively reduces errors introduced during KV Cache quantization.
Group-Wise Low-Bit Quantization. To minimize the memory usage of the quantized KV Cache, we apply group-wise lowest possible bit quantization for KV. The quantization process is detailed as follows: ① Calculation of scaling factor. The input elements are divided into groups based on the asymmetric quantization strategy (e.g., per-channel for Key or per-token for Value). For each group, we first determine the minimum ($min\_val$) and maximum ($max\_val$) values. The scaling factor $s$ is then calculated as: $s=\frac{max\_val-min\_val}{q_{max}}$, where $q_{max}$ denotes the maximum quantized value. ② Element quantization. Each element is quantized using: $q=round\left(\frac{x-min\_val}{s}\right)$, where $x$ represents the original element value, and $q$ is the quantized value. ③ Clipping of quantized values. The quantized values are constrained to the range $[0,q_{max}]$ through clipping: $q=\max(0,\min(q,{q_{max}}))$. ④ Storage and dequantization. The quantized values are stored using bit operations within a 32-bit integer ($int32$). For 4-bit, 2-bit, and 1-bit quantization, the number of elements in per $int32$ is: $feat\_per\_int=32/bit$. During inference, dequantization is performed using: $x=q\cdot s+min\_val$. For 3-bit quantization, we introduce a new packing strategy to maximize memory efficiency. We organize the quantized elements into blocks of 11, each stored in a 32-bit integer, with the first 10 elements quantized to 3 bits and the 11th element to 2 bits. The clipping range is adjusted based on the element index:
$q_{max}=\begin{cases}7,&i=0,1,...,9\\ 3,&i=10\end{cases}$ (3)
where $i$ is the element index within a block. This method provides a 10% increase in packing density compared to a uniform 3-bit quantization that can only hold 10 elements per $int32$.
3.3 Dynamic Pivotal Context Selection
In the KV Cache, not all Keys and Values are equally important for generating future tokens. Recent tokens provide the most relevant contextual information for the generation of subsequent tokens, and usually have more impact on the upcoming tokens. We define the Keys and Values corresponding to these pivotal recent tokens as the Recent Pivotal Context (RPC). To optimize model performance while minimizing memory usage, we propose a dynamic RPC selection strategy based on the importance analysis provided by KVmix profiler. Specifically, for the $i$-layer, we assign it an RPC selection ratio $r$ based on the $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ scores, with higher $\bar{s}_{k_{i}}$ and $\bar{s}_{v_{i}}$ resulting in larger $r$. The number of RPCs is computed by $num\_RPC=\lfloor r\times current\_RPC\rfloor$. $current\_RPC$ is the sum of the number of new Key and Value states at current time step and the number of historical RPCs.
Figure 3: Dynamic adjustment process of KV Cache quantization in the prefill and decoder phases.
The corresponding number of KV pairs is selected as RPCs based on $num\_RPC$. We keep full precision for the RPCs while performing the mixed quantization for the less critical and older KV pairs. This strategy ensures that the number of full-precision RPCs is dynamically reduced in real-time during long context inference, thus avoiding excessive memory pressure caused by preserving a large number of full-precision KV pairs, while maintaining high-quality sequence generation. In addition, since the importance of Keys and Values may differ within the same layer, the RPC selection ratio for Keys and Values varies within that layer. The RPC selection ratio can be adjusted to provide flexibility in the dynamic balance between inference accuracy and memory usage.
3.4 CUDA Implementation
During model inference, the quantization of the KV Cache introduces additional overhead from quantization and dequantization operations. To mitigate this overhead and enhance inference efficiency, we design efficient CUDA kernels for quantization, dequantization, and matrix-vector multiplication. ① Fusion of quantization and concatenation in the decoding phase. In the decoding phase, the Key states and Value states of the current layer need to be concatenated with the historical KV Cache. Directly quantizing the current states before concatenation incurs additional memory access overhead. Therefore, we fuse the quantization process with the concatenation operation into a single CUDA kernel. Specifically, the kernel processes each element in a streaming manner: each element of the Key states and Value states is quantized and immediately appended to the quantized historical KV Cache, which reduces memory access overhead. We also leverage CUDA thread blocks to process data from different tokens in parallel and utilize shared memory to cache intermediate results, further improving data locality. ② Fusion of dequantization and matrix-vector multiplication in attention computation. During attention computation, the quantized KV Cache must be dequantized before performing matrix-vector multiplication. Dequantizing the entire KV Cache prior to multiplication increases memory usage. To mitigate this issue, we fuse dequantization with matrix-vector multiplication. Specifically, each element is dequantized and immediately multiplied and accumulated with its corresponding element, minimizing additional memory overhead. ③ Efficient kernels for multi-bit quantization configurations. To support the various quantization bit-widths of KVmix, we develop efficient CUDA kernels for 1-bit, 2-bit, 3-bit, and 4-bit quantization. Additionally, we implement customized matrix-vector multiplication kernels for each quantization configuration, ensuring compatibility with varying bit-widths.
4 Experimental Results
4.1 Experimental Setup
We evaluated the proposed method using the Llama 2-7B-hf [32], Llama 3-8b-Instruct, Llama 3.1-8B [33], and Mistral-7B-Instruct-v0.3 [34] models. The datasets were selected based on three distinct evaluation schemes: ① Long Context Evaluation: We used the LongBench [35] benchmark to assess performance on long-context tasks. This benchmark encompasses key long-text application scenarios, including single-document question answering (QA), multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The maximum sequence length for all tasks was set to 4096. ② Language Modeling: We measured the perplexity of our method on the Wikitext-2 [36] dataset to evaluate its language modeling capabilities. ③ Mathematical Reasoning: We employed the GSM8K [37] dataset to assess the model’s performance on mathematical reasoning tasks.
We used the NVIDIA RTX 4090 GPU (24GB) to evaluate the quantized model’s inference efficiency and the KV cache’s compression rate. Section 4.2 elaborates on the specific configurations for KV quantization bits and the proportion of RPC.
4.2 Profiling Results
We used 3-bit and 2-bit mixed quantization for Key, and 4-bit and 2-bit mixed quantization for Value. The proportion of 4-bit and 3-bit is 20% of the model layers, and the remaining model layers are all quantized with 2-bit, that is, the average quantization bit of Key is 2.19 (exact value is 2.1875), and the average quantization bit of Value is 2.38 (exact value is 2.375). We randomly selected 30 prompts from the LongBench for KV importance analysis. Using the KVmix profiler, we obtained the KV bit configurations and RPC proportions for each layer. Fig. 4 shows the detailed configurations. When Key and Value are quantized to 3 bits and 4 bits, their RPC proportions are set to 20%; for 2-bit quantization, the RPC proportions are set to 10%. The group size for quantization is 32. The following experiments use this configuration unless otherwise specified. Users can flexibly customize the quantization configuration by modifying the proportion of layers with different bits and the proportion of RPC in KVmix profiler according to different model accuracy or memory requirements.
Figure 4: Detailed quantization configurations of KVmix-k2.19v2.38 for different models.
4.3 Performance Evaluation
4.3.1 Long Context
We evaluated the performance of various quantization configurations across 8 distinct datasets from the LongBench benchmark, using the FP16 model as the baseline. The detailed experimental results are presented in the Table 1. The findings indicate that KVmix-k2.19v2.38 achieves an average accuracy loss (average values on 4 different models) of 0.92% compared to the FP16 baseline. In contrast, the average accuracy loss of KVmix-2bit reached 2.14% relative to the baseline. In addition, random selection of high-bit quantization layers (random-k2.19v2.38) leads to an average accuracy loss of 1.78% compared to the baseline. These accuracy losses are significantly higher than KVmix-k2.19v2.38, demonstrating the advantage of KV importance-aware mixed quantization. Appendix C presents additional evaluation results on LongBench. Increasing the proportion of high-bit quantization layers for KV to 30% (KVmix-k2.28v2.56) significantly improves model accuracy compared to KVmix-k2.19v2.38, nearly matching the performance of full 4-bit quantization. Certainly, compared to KVmix-k2.19v2.38, KVmix-k2.28v2.56 increases memory usage.
Table 1: Model accuracy of 4 LLMs on LongBench with different quantization configurations. KVmix-k2.19v2.38 uses the configurations of Section 4.2. KVmix-2bit uses the asymmetric 2-bit (Key per-channel and Value per-token) quantization for all model layers (RPC ratio is set to 10%). random-k2.19v2.38 randomly selects 20% of the model layers to perform asymmetric 3-bit and 4-bit quantization for Key and Value (RPC ratio is set to 20%), and the remaining layers are 2-bit quantization (RPC ratio is set to 10%).
Models Methods Datasets Average
TriviaQA Qasper MF-en QMSum 2WikiMQA Rbench-P TREC PsgRetr-en
Llama-2-7B-hf FP16 78.89 9.55 22.86 21.19 9.94 55.64 66.00 6.64 33.839
KVmix-2bit 77.57 9.58 22.47 20.45 9.15 56.34 66.00 5.29 33.356
random-k2.19v2.38 78.30 9.39 22.54 20.41 9.46 56.36 66.00 5.49 33.494
KVmix-k2.19v2.38 78.78 9.59 22.82 20.49 9.77 56.54 66.00 5.72 33.714
Llama-3-8B-Instruct FP16 78.35 40.75 46.80 21.69 32.39 49.77 70.50 37.00 47.156
KVmix-2bit 76.13 39.18 45.70 21.20 32.19 44.56 71.00 36.30 45.783
random-k2.19v2.38 78.01 39.17 45.90 21.22 32.02 45.36 71.00 36.50 46.148
KVmix-k2.19v2.38 78.13 39.15 46.31 21.26 32.20 47.56 71.00 36.50 46.514
Llama-3.1-8B FP16 83.67 11.53 31.13 22.88 13.92 61.84 67.50 19.50 38.996
KVmix-2bit 83.10 10.90 30.76 22.11 13.08 58.92 67.00 19.00 38.109
random-k2.19v2.38 83.25 10.90 31.05 22.34 13.05 59.26 67.00 19.00 38.231
KVmix-k2.19v2.38 83.28 11.40 31.49 22.90 12.92 59.96 67.50 19.50 38.619
Mistral-7B-Instruct-v0.3 FP16 84.29 36.19 54.70 21.79 35.08 53.06 73.50 32.50 48.889
KVmix-2bit 84.08 34.29 53.87 21.37 33.39 50.99 73.50 32.00 47.936
random-k2.19v2.38 84.01 34.35 53.61 21.45 33.40 50.59 73.50 32.50 47.926
KVmix-k2.19v2.38 84.03 35.67 53.68 21.84 33.81 51.98 73.50 32.75 48.408
4.3.2 GSM8K and Wikitext-2
We evaluated the capabilities of the quantized model in language modeling and mathematical reasoning using the FP16 model as a baseline. The evaluation was conducted using the lm_eval [38] framework, where the quantized model replaced the Hugging Face model. Specifically, we measured the accuracy on the GSM8K dataset and the perplexity on the Wikitext-2 dataset. The experimental results are detailed in Table 2. The results show that 2bit (k-T, v-T) suffers a catastrophic performance loss on GSM8K and Wikitext-2, and the model almost loses its reasoning ability. For 4bit (k-T, v-T), the performance loss of the model on GSM8K and Wikitext-2 also reached 9.17% and 5.28%, respectively.
Table 2: Model accuracy on GSM8K and perplexity on Wikitext-2 using Llama 2-7B-hf. 2bit (k-T, v-T) uses the symmetric 2-bit (Key per-token and Value per-token) quantization for all model layers, and 4bit (k-T, v-T) uses the symmetric 4-bit quantization; their RPC ratio is set to 0.
Methods GSM8K ($\uparrow$) Wikitext-2 ($\downarrow$)
FP16 13.52 8.71
2bit (k-T, v-T) 0.83 11089
4bit (k-T, v-T) 12.28 9.17
KVmix-2bit 11.80 8.73
random-k2.19v2.38 11.97 8.72
KVmix-k2.19v2.38 13.25 8.71
In contrast, on the Wikitext-2, the perplexity score of KVmix-k2.19v2.38 is almost comparable to the baseline, while on the more challenging GSM8K mathematical reasoning task, KVmix-k2.19v2.38 has an accuracy loss of 2.00%, which significantly outperforms the 2bit (k-T, v-T) and the 4bit (k-T, v-T). Moreover, on GSM8K, KVmix-k2.19v2.38 shows a significant accuracy improvement compared to KVmix-2bit and random-k2.19v2.38, which do not leverage the KV importance analysis for more accurate quantization. These results demonstrate the superior performance of KVmix-k2.19v2.38 in complex task reasoning.
4.4 Inference Efficiency and Memory Usage Evaluation
We evaluated the inference throughput and memory usage of KVmix during inference on an NVIDIA RTX 4090 GPU (24GB). To ensure a fair comparison, we applied identical input data across all evaluated methods. The number of input tokens is 688, the maximum number of new tokens is set to 1024, and the model is Llama 2-7B-hf. We compared KVmix against the methods optimized for memory and throughput: KIVI [13] with 2-bit quantization (full precision residual of 64) and Atom [23] with 4-bit quantization. Memory usage results are illustrated in Fig. 6, with a batch size fixed at 4. The reported memory usage represents the dynamic peak memory consumption during inference minus the memory occupied by the model before inference. To fully utilize the GPU memory, we incrementally increased the batch size to explore KVmix’s maximum inference throughput. The throughput results are shown in Fig. 6. The baseline (FP16), Aotm-4bit, and KIVI-2bit-r64 reach out of memory at batch sizes of 4, 18, and 28, respectively, while the KVmix-k2.19v2.38 can reach a maximum batch size of 30 with an inference throughput of 1032 tokens per second.
The results reveal that KVmix-k2.19v2.38 achieves a 4.9$\times$ reduction in inference runtime memory usage and up to a 5.3$\times$ increase in throughput compared to the baseline. This efficient memory compression stems from KVmix’s extremely low bit quantization and dynamic RPC strategy, which progressively reduces the full-precision KV Cache as inference progresses. In contrast, KIVI employs a fixed full-precision residual strategy, unable to dynamically reduce the number of full-precision KVs. Thus, KVmix saves more memory than KIVI’s 2-bit quantization despite using Key-2.19 and Value-2.38 bit configuration. Meanwhile, Atom-4bit quantizes both model weights and activations while utilizing Tensor Cores for optimized kernel execution, achieving a higher throughput at the same batch size but incurring greater model accuracy degradation, for example, it has an average accuracy loss of nearly 10% using the Llama 2-7B-hf model on the Wikitext-2 dataset.
Figure 5: Dynamic peak memory usage of different methods during inference on the Llama 2-7B-hf model with a batch size of 4.
Figure 6: Inference throughput of different quantization methods with different batch sizes on the Llama 2-7B-hf model.
4.5 Comparison with SOTA Methods
We compared KVmix against prior SOTA methods for KV Cache, specifically Key per-channel and Value per-token methods, i.e., KIVI [13] and KVQuant [22], since they are known to minimize KV quantization errors. Table 3 presents the model accuracy comparison on the LongBench. The results show that KVmix-k2.19v2.38 outperforms KIVI-2bit-r64, and reduces the average accuracy loss by 1.57%. Although KVQuant-3bit-1% is competitive with KVmix-k2.19v2.38 in terms of accuracy, it achieves at most 4.8$\times$ memory compression and 1.70$\times$ inference acceleration using Llama 2-7B-hf, which is lower than the 4.9$\times$ memory compression and 5.32$\times$ inference acceleration achieved by KVmix-k2.19v2.38. By increasing the quantization bits of KVmix, the advantage of KVmix-k2.28v2.56 over KVQuant in terms of model accuracy is now obvious. Moreover, KVmix-k2.28v2.56 maintains a 4.8$\times$ memory compression and a 5.23$\times$ inference acceleration. This flexibility in balancing accuracy and quantization bits is also one of the important advantages of KVmix.
Table 3: Accuracy comparison of different quantization methods on LongBench using the Llama 2-7B-hf model. KIVI-2bit-r64 uses 2-bit quantization with a full-precision residual of 64. KVQuant-3bit-1% uses 3-bit quantization and 1% outlier handling. KVmix-k2.28v2.56 increases the proportion of high-bit quantization layers in KVmix-k2.19v2.38 to 30%, its detailed configurations are shown in Appendix C, Fig. 8.
Methods TriviaQA Qasper MF-en QMSum 2WikiMQA Repobench-P TREC PsgRetr-en Average
FP16 78.89 9.55 22.86 21.19 9.94 55.64 66.00 6.64 33.839
KIVI-2bit-r64 77.08 9.16 22.55 20.12 9.05 56.15 66.00 5.62 33.193
KVQuant-3bit-1% 78.79 10.51 22.61 20.58 9.75 55.62 66.00 5.76 33.703
KVmix-k2.19v2.38 78.78 9.59 22.82 20.49 9.77 56.54 66.00 5.72 33.714
KVmix-k2.28v2.56 78.05 10.21 23.21 20.63 9.72 56.61 66.00 6.08 33.814
5 Conclusion
This paper addresses the memory bottleneck issue posed by the KV Cache during the inference of LLMs by proposing a novel mixed quantization method, KVmix. KVmix achieves layer importance-aware mixed-precision quantization by leveraging KV weight gradients. It also provides dynamic long-context optimization to reduce memory usage while preserving generation quality. Experiments show that KVmix, with an average of 2.19 bits for Keys and 2.38 bits for Values, reduces inference memory usage by 4.9$\times$ and increases throughput by 5.3$\times$, while maintaining near-lossless inference quality. KVmix offers flexibility and ease of use, enabling the adjustment of quantization strategies based on specific application scenarios, thus effectively balancing model accuracy, inference efficiency, and memory usage. However, it has not yet achieved real-time mixed quantization during inference. Future work will explore lightweight mechanisms for real-time bit adjustments in mixed-precision quantization, enabling KVmix to dynamically respond to changing inference demands while maintaining efficiency.
Appendix A KV Weight Analysis
Fig. 7 presents the heatmaps of the KV weights ($W_{k}$ and $W_{v}$) for the Llama 3-8B-Instruct and Mistral-7B-Instruct-v0.3 models. For different layers of various models, the $W_{k}$ and $W_{v}$ exhibit significant differences. The differences are consistent with the conclusions in Section 2.2.
Figure 7: Projection matrix weights of K and V across different model layers for Llama 3-8b-Instruct (top) and Mistral-7B-Instruct-v0.3 (bottom). The heatmap visualizes the projection weights by sampling every sixth element in the matrices. Additionally, the weight values along the axes have been normalized. "Norm" represents the L2 norm of the weight matrix for each layer, while "Range" indicates the range of values within each layer’s weight matrix.
Appendix B KVmix Profiler Algorithm
In the Algorithm B, we provide the workflow of the KVmix profiler, including the process from data processing, KV importance analysis, and quantization bit configuration. When using the KVmix profiler, selecting 20 to 30 prompts is sufficient to yield reliable importance analysis results. It typically takes only 10 to 15 minutes to complete on a single NVIDIA RTX 4090 GPU.
  Algorithm 1 Model Profiling and Quantization
Model $model$, inputs $input\_ids$, $attention\_mask$, $labels$, layers $num\_layers$, datasets $datasets$
Quantization bits $quant\_bits$
for dataset in datasets do
     prompts $\leftarrow$ randomSample(dataset, n);  $\triangleright$ Sample n prompts
     inputs $\leftarrow$ tokenize(prompts);  $\triangleright$ Prepare inputs
     model $\leftarrow$ loadFullModel();  $\triangleright$ Load full-precision model
     kScores, vScores $\leftarrow$ calcImportance(model, inputs, num_layers);  $\triangleright$ Compute importance
     kLayers, vLayers $\leftarrow$ classify(kScores, vScores, num_layers);  $\triangleright$ Classify layers
     quantBits $\leftarrow$ setBits(kLayers, vLayers);  $\triangleright$ Set quantization bits
     del model;  $\triangleright$ Free memory
     model $\leftarrow$ loadQuantModel(quantBits);  $\triangleright$ Load quantized model
     preds $\leftarrow$ infer(model, dataset);  $\triangleright$ Inference
     save(preds);  $\triangleright$ Save results
     del model;  $\triangleright$ Free memory
end for
Function calcImportance(model, inputs, num_layers):  $\triangleright$ Compute importance scores
kScoresAll, vScoresAll $\leftarrow$ [], [];
for input in inputs do
     kScores, vScores $\leftarrow$ [], [];
     for layer in range(num_layersdo
         loss $\leftarrow$ model.forward(input);  $\triangleright$ Compute loss
         kGrad $\leftarrow$ grad(loss, kParams(layer));  $\triangleright$ Gradient for Key
         vGrad $\leftarrow$ grad(loss, vParams(layer));  $\triangleright$ Gradient for Value
         kScores.append(kGrad.norm());  $\triangleright$ Append Key score
         vScores.append(vGrad.norm());  $\triangleright$ Append Value score
         clearMemory();  $\triangleright$ Free memory
     end for
     kScoresAll.append(kScores);
     vScoresAll.append(vScores);
end for
kScoresMean $\leftarrow$ mean(kScoresAll);  $\triangleright$ Average Key scores
vScoresMean $\leftarrow$ mean(vScoresAll);  $\triangleright$ Average Value scores
return kScoresMean, vScoresMean;
Appendix C Additional Performance Evaluation of KVmix with LongBench
We present extended experimental results building upon the experiments detailed in Section 4.3.1, aiming to provide a more comprehensive performance comparison and to illustrate the impact of different KVmix configurations on model accuracy. Fig. 8 displays the detailed configuration results when the proportion of high-bit quantization layers is increased to 30%, corresponding to KVmix-k2.28v2.56. Table 4 summarizes the performance comparison of different models using KVmix across various quantization configurations. The experimental results demonstrate that KVmix-k2.19v2.38 achieves a significant accuracy improvement over KVmix-2bit and random-k2.19v2.38. Compared to KVmix-4bit, KVmix-k2.19v2.38 has an average accuracy loss of within 0.013, but achieves nearly twice the KV Cache compression. Further increasing the proportion of high-bit quantization layers to 30% (KVmix-k2.28v2.56) leads to a noticeable improvement in model accuracy, almost the same level as KVmix-4bit, validating the effectiveness and flexibility of KVmix.
Figure 8: Detailed quantization configuration of KVmix-k2.28v2.56 in different models.
Table 4: Model accuracy of 4 LLMs on LongBench with different quantization configurations. KVmix-k2.19v2.38 uses the configurations of Section 4.2. KVmix-k2.28v2.56 uses the configurations of Fig. 8. KVmix-2bit uses the asymmetric 2-bit (Key per-channel and Value per-token) quantization for all model layers (RPC ratio is set to 10%). KVmix-4bit uses the asymmetric 4-bit quantization for all model layers (RPC ratio is set to 20%). random-k2.19v2.38 randomly selects 20% of the model layers to perform asymmetric 3-bit and 4-bit quantization for Key and Value (RPC ratio is set to 20%), and the remaining layers are 2-bit quantization (RPC ratio is set to 10%).
Models Methods Datasets Average
TriviaQA
Qasper
MF-en
QMSum
2WikiMQA
GovRep
Rbench-P
LCC
TREC
PsgRetr-en
Llama-2-7B-hf FP16 78.89 9.55 22.86 21.19 9.94 17.36 55.64 66.70 66.00 6.64 35.477
KVmix-4bit 77.88 9.11 22.43 21.13 9.75 17.53 56.17 66.70 66.00 6.35 35.305
KVmix-2bit 77.57 9.58 22.47 20.45 9.15 14.03 56.34 66.18 66.00 5.29 34.706
random-k2.19v2.38 78.30 9.39 22.54 20.41 9.46 14.01 56.36 66.31 66.00 5.49 34.827
KVmix-k2.19v2.38 78.78 9.59 22.82 20.49 9.77 15.45 56.54 66.59 66.00 5.72 35.175
KVmix-k2.28v2.56 78.05 10.21 23.21 20.63 9.72 15.62 56.61 66.64 66.00 6.08 35.277
Llama-3-8B-Instruct FP16 78.35 40.75 46.80 21.69 32.39 30.62 49.77 56.51 70.50 37.00 46.438
KVmix-4bit 78.30 41.39 46.31 21.58 32.80 30.31 49.21 56.73 70.50 37.00 46.413
KVmix-2bit 76.13 39.18 45.70 21.20 32.19 29.62 44.56 49.10 71.00 36.30 44.498
random-k2.19v2.38 78.01 39.17 45.90 21.22 32.02 29.55 45.36 49.71 71.00 36.50 44.844
KVmix-k2.19v2.38 78.13 39.15 46.31 21.26 32.20 29.95 47.56 51.62 71.00 36.50 45.368
KVmix-k2.28v2.56 78.27 39.30 46.42 21.11 32.82 29.92 49.13 53.92 71.00 37.00 45.889
Llama-3.1-8B FP16 83.67 11.53 31.13 22.88 13.92 29.23 61.84 68.96 67.50 19.50 41.016
KVmix-4bit 83.77 11.75 31.44 22.59 13.85 29.04 61.20 69.28 67.50 19.00 40.942
KVmix-2bit 83.10 10.90 30.76 22.11 13.08 27.02 58.92 68.13 67.00 19.00 40.002
random-k2.19v2.38 83.25 10.90 31.05 22.34 13.05 27.01 59.26 68.39 67.00 19.00 40.125
KVmix-k2.19v2.38 83.28 11.40 31.49 22.90 12.92 27.22 59.96 68.54 67.50 19.50 40.471
KVmix-k2.28v2.56 83.80 11.37 31.31 22.85 13.47 27.75 60.94 68.10 67.50 19.50 40.659
Mistral-7B-Instruct-v0.3 FP16 84.29 36.19 54.70 21.79 35.08 32.84 53.06 57.56 73.50 32.50 48.151
KVmix-4bit 84.49 35.83 54.35 22.39 35.08 32.83 52.69 57.99 73.50 32.50 48.165
KVmix-2bit 84.08 34.29 53.87 21.37 33.39 32.05 50.99 56.18 73.50 32.00 47.172
random-k2.19v2.38 84.01 34.35 53.61 21.45 33.40 32.20 50.59 56.53 73.50 32.50 47.214
KVmix-k2.19v2.38 84.03 35.67 53.68 21.84 33.81 32.19 51.98 56.88 73.50 32.75 47.633
KVmix-k2.28v2.56 84.45 35.87 54.35 21.97 34.23 32.51 51.85 57.25 73.50 33.00 47.898